Fermi National Laboratory


Accelerator Update - June 4-June 6

Wednesday June 4
The day shift began with Operations monitoring store 2636, with the store 2642, with the store 2636, with the Antiproton Source (Pbar) store 2636, with the stacking, and with MiniBooNE taking beam.

At 11:05 AM, Operations terminated store 2642.

The Collider Detector Facility (CDF) and D-zero (D0) experimenters store 2636, with the accessed their collision halls at 11:26 AM. During this access, store 2636, with the Tevatron (TeV) technicians worked on three power supplies.

Both CDF and D0 experiments were out of access by 1 PM.

store 2636, with the Switchyard experts completed a study period at 1:27 PM.

Operations put the TeV into store 2636, with the shot setup at 2:14 PM.

At 3:23 PM, Operations rested a store 2636, with the Linac store 2636, with the RF station (LRF1), in an attempt to extend its life.

A Pbar RF station (ARF1) tripped off at 3:25 PM, delaying the shots of antiprotons. This station reset okay, but then another station (ARF4) took over the delaying action. An expert fixed the problem and Operations loaded the final antiprotons at 3:58 PM.

At 4:22 PM, Operations established store 2644 with an initial store 2636, with the luminosity of 28.72E30. Operators turned on LRF1. Pbar resumed stacking.

A TeV RF station (TRF5) tripped off at 6:04 PM on a heater ground fault indication. Per expert instructions, operators reset the station at 8:13 PM.

At 6:12 PM, Operations resumed sending beam to MiniBooNE.

Unfortunately, a TeV store 2636, with the sector B4 store 2636, with the quench caused the store to store 2636, with the abort at 6:28 PM. Cryo experts discovered a remotely controlled store 2636, with the cryo system valve had opened twenty-five percent more that its normal position. They made a temporary fix to the control loops until the software cause of the problem can be found.

D0 accessed their collision hall at 7:20 PM. They came out of access two hours later.

TeV experts began an store 2636, with the aperture scan study at 11:06 PM.

Thursday June 5
The midnight shift began with Pbar stacking, with TeV experts conducting studies, and with MiniBooNE taking beam.

At 3:17 AM, a store 2636, with the Booster RF station (BRF13) tripped off for the first of four times over a forty-five minute period.

Operations put the TeV into shot setup at 5:08 AM.

At 5:50 AM, operators discovered some devices in the AP2 and AP3 transfer lines without power. A breaker had tripped off. The Duty Electrician had the breaker reset by 6:19 AM.

Operation established store 2646 at 7:29 AM. Its initial luminosity was 26.53E30. Pbar resumed stacking, but beam was not sent to MiniBooNE.

At 8:57 AM, the Linac beam was turned off so techs could change out the power amplifier tube for LRF1.

Linac techs completed their work on LRF1 and at 12:30 PM Operations established beam to the Linac dump. Pbar resumed stacking. Fifteen minutes later, MiniBooNE resumed taking beam.

At 2:17 PM, a poor Linac tune was preventing MiniBooNE from getting beam. Operators found a problem with a quadrupole power supply readback. After an hour of troubleshooting, Operations reestablished beam to MiniBooNE.

Operations found a Booster kicker (MKS04) that wasn't working at 4:39 PM. A Booster expert investigated the problem. At 6 PM, he reported that the problem was in the tunnel.

At 5:45 PM, two store 2636, with the Debuncher vacuum valves closed due to vacuum pumps tripping off. Operations reset the pumps. Despite all efforts, the vacuum eventually got worse and experts were called in.

A Debuncher store 2636, with the septum (D:ESEPV) tripped off at 6:02 PM. It reset okay.

At 7:28 PM, Operations killed the stack and prepared Pbar and Booster for access.

The vacuum techs investigating the Pbar vacuum problem reported a store 2636, with the LCW leak in sector AP10 at 7:41 PM. Water techs came in and repaired a bad hose (on A1S12).

A Booster expert began repairs on MKS04 at 8:55 PM.

Vacuum techs repaired a leaking flange. They were out of the Pbar tunnel by 10:31 PM.

The Booster expert completed his work at 10:41 PM.

Operations established beam to the Booster by 11:07 PM, established beam to MiniBooNE by 11:19 PM, and reestablished Pbar stacking by 11:30 PM.

Friday June 6
The midnight shift began with Operations monitoring store 2646, with Pbar stacking, with MiniBooNE taking beam, and with Recycler experts conducting studies.

At 6:04 AM, operations terminated store 2646. TeV experts began TeV ramping studies ten minutes later.

Today's Plan
The plan for today is to let TeV experts continue with their studies until noon and then put the TeV into shot setup. The plan for the weekend is to stack and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/6/2003   email Fermilab